# Risk-Aware Adaptive Control Barrier Functions for Safe Control of Nonlinear Systems under Stochastic Uncertainty

### v2

@@ II-B Coherent Risk Measure
[…]
@@
The one-step conditional risk measures $\rho_{t}:\mathcal{H}_{t+1}\to\mathcal{H}_{t}$, $t=1,\dots,N-1$, as in (9), are called coherent risk measures if they satisfy the following properties:
@@
  • Translational Invariance: $\rho_{t}(h+c)=c+\rho_{t}(h)$, for all $h\in\mathcal{H}_{t-1}$ and $c\in\mathbb{R}$.
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
Let $\varphi_{i,j}(\bm{\pi}_{i}),\ j\in\{1,...,m+1-i\}$ and $\mathcal{B}_{i,j},\ j\in\{0,...,m-i\}$ be defined by (17) and (18) respectively. By Def. 2, a function $h_{i}:\mathbb{R}^{m+1-i}\to\mathbb{R}$ is a DHOCBF with relative degree $m+1-i$ for system (16) if there exist class $\kappa$ functions $\alpha_{i,j},\ j\in\{1,...,m+1-i\}$ as in (17) such that
@@
For simplicity, we can make $h_{i}(\bm{\pi}_{i,t})=\pi_{i,1,t}=a_{i,t}.$ Based on Rem. 1, each $a_{i,t}$ will be positive. The remaining question is how to define an adaptive DHOCBF to guarantee $h(\mathbf{x})\geq 0$ with the assistance of auxiliary variables. Let $\bm{\Pi}_{t}\coloneqq[\bm{\pi}_{1,t},\dots,\bm{\pi}_{m,t}]^{T},\bm{\nu}_{t}\coloneqq[\nu_{1,t},\dots,\nu_{m,t}]^{T}$ denote the auxiliary states and control inputs of system (16). We can define a sequence of functions as follows: We can define a sequence of functions
@@
where $i\in\{0,...,m-1\}.$ Since $a_{i,t}$ is a DHOCBF with relative degree $m+1-i$ for (16), based on (19), we define a constraint set $\mathcal{U}_{\bm{a}}$ for $\bm{\nu}_{t}$ as follows:
@@
Let $\psi_{i}(\mathbf{x},\bm{\Pi}),\ i\in\{1,...,m\}$ be defined by (22) and $\mathcal{C}_{i},\ i\in\{0,...,m-1\}$ be defined by (23). A function $h(\mathbf{x}):\mathbb{R}^{n}\to\mathbb{R}$ is a Discrete-Time Auxiliary-Variable Adaptive Control Barrier Function (DAVCBF) with relative degree $m$ for system (1) if every $a_{i,t},i\in\{1,...,m\}$ is a DHOCBF with relative degree $m+1-i$ for the auxiliary system (16), and there exist class $\kappa$ functions $\alpha_{j},j\in\{1,...,m\}$ s.t.
@@ IV-B Risk Sensitive Forward Invariance
$\rho_{0,t}(0,0,\dots,\psi_{i}(\mathbf{x}_{t},\bm{\Pi}_{t}))\geq 0,\quad\forall t\in\mathbb{N}_{\geq 0}.$ (26)
@@ IV-C Ensuring Risk Sensitive Safety with RACBFs
Let $\psi_{i}(\mathbf{x},\bm{\Pi}),\ i\in\{1,...,m\}$ be defined by (22), $\mathcal{C}_{i},\ i\in\{0,...,m-1\}$ be defined by (23) and a dynamic coherent risk measure $\rho$ be defined by Def. 4. A function $h(\mathbf{x}):\mathbb{R}^{n}\to\mathbb{R}$ is a Risk-Aware Adaptive Control Barrier Function (RACBF) with relative degree $m$ for system (13) if every $a_{i,t},i\in\{1,...,m\}$ is a DHOCBF with relative degree $m+1-i$ for the auxiliary system (16), and there exist class $\kappa$ functions $\alpha_{j},j\in\{1,...,m\}$ for $\psi_{i}(\mathbf{x},\bm{\Pi})$ s.t.
@@ IV-D Optimal Control with RACBFs
$\begin{split}-(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})\geq 0,\sigma_{i,t}\geq 0,\\ \sigma_{i,t}\geq-\rho^{i}(\psi_{m}(\mathbf{x}_{t},\bm{\Pi}_{t}))-\zeta_{t},i\in\{1,...,|\mathcal{W}|\}.\end{split}$ (31)
@@
where $\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})=\frac{\psi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{i,t}},\alpha_{i}(\psi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t}))=\lambda_{i}a_{i,t}\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t}),0<\lambda_{i}\leq 1,i\in\{1,\dots,m\}$. The term $\frac{a_{i,t}}{a_{i,t+1}}$ can be adjusted adaptable to ameliorate the conservativeness of control strategy that $(1-\lambda_{i})\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})$ may have, which allows for a broader range of feasible control inputs and confirms the adaptivity of DAVCBFs to control constraint and conservativeness of control strategy. The aforementioned benefits can be further amplified by formulating the optimal control problem in the model predictive control framework.
@@
Compared to Eqn. (32), we observe that $(1-\lambda_{i}a_{i,t})$ has a range of $[-\infty,1]$, while $\frac{a_{i,t}}{a_{i,t+1}}(1-\lambda_{i})\phi_{i-1}$ ranges from $[0,+\infty]$. The form of Eqn. (32) maximizes the flexibility in adjusting the parameters of the class $\kappa$ function while eliminating negative parameters that may compromise safety.
@@ V CASE STUDY AND SIMULATIONS
where $M$ denotes the mass, and $\mathbf{x}_{t}=[x_{t},y_{t},\theta_{t},v_{t}]^{T}$ captures the 2-D location, heading angle, and linear speed; $\mathbf{u}_{t}=[u_{1,t},u_{2,t}]^{T}$ represents angular velocity ($u_{1,t}$) and driven force ($u_{2,t}$), respectively. The disturbance $\mathbf{w}_{t}\in\mathcal{W}$ enters the system linearly and is characterized by a probability mass function over the states. In this study, the probability mass function is a simple normal distribution centered at 00 with a standard deviation of $[0.2,0.2,0.2,0.2]$ for the four states $[x_{t},y_{t},\theta_{t},v_{t}]$. The system is discretized with $\Delta t=0.1s$ and the total number of discretization steps $t_{T}$ in Problem. 1 is 50. System (34) is subject to the following input constraint:
@@
The initial state is $[-1.5,0,\frac{\pi}{12},1]^{T}$ and the target state is $\mathbf{x}_{e}=[2,0,0,0]^{T},$ which are marked as black diamond and green circle in Fig. 1. The circular obstacle is centered at $(x_{o},y_{o})=(0,0)$ with $r_{o}=1$, which is displayed in pale red. To initialize the proposed RACBFs, we multiply an auxiliary variable $a_{1,t}$ by a quadratic distance function for circular obstacle avoidance as $\psi_{0}(\mathbf{x}_{t},\bm{{\pi}}_{1,t})=a_{1,t}((x_{t}-x_{o})^{2}+(y_{t}-y_{o})^{2}-r_{o}^{2})$. Since the relative degree of $\psi_{0}$ with respect to system (34) is 2, the DHOCBFs for $a_{1,t}$ are defined as

### v3

@@ II-B Coherent Risk Measure
$\begin{split}\rho_{t:N}(h_{t},\dots,h_{N})=\rho_{t}(h_{t}+\rho_{t+1}(h_{t+1}+\rho_{t+2}(h_{t+2}+\dots{\\ }+\rho_{N-1}(h_{N-1}+\rho_{N}(h_{N}))\dots))).\end{split}$ (9)
@@
The one-step conditional risk measures $\rho_{t}:\mathcal{H}_{t}\to\mathcal{H}_{t-1}$, $t=1,\dots,N$, as in (9), are called coherent risk measures if they satisfy the following properties:
@@
  • Translational Invariance: $\rho_{t}(h+c)=c+\rho_{t}(h)$, for all $h\in\mathcal{H}_{t}$ and $c\in\mathbb{R}$.
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
Let $\varphi_{i,j}(\bm{\pi}_{i}),\ j\in\{1,...,m+1-i\}$ and $\mathcal{B}_{i,j},\ j\in\{0,...,m-i\}$ be defined by (17) and (18) respectively. By Def. 2, a function $h_{i}:\mathbb{R}^{m+1-i}\to\mathbb{R}$ is a DHOCBF with relative degree m+1im+1-i for system (16) if there exist class $\kappa$ functions $\alpha_{i,j},\ j\in\{1,...,m+1-i\}$ as in (17) such that
@@
For simplicity, we can make $h_{i}(\bm{\pi}_{i,t})=\pi_{i,1,t}=a_{i,t}.$ Based on Rem. 1, each $a_{i,t}$ will be positive. The remaining question is how to define an adaptive DHOCBF to guarantee $h(\mathbf{x})\geq 0$ with the assistance of auxiliary variables. Let $\bm{\Pi}_{t}\coloneqq[\bm{\pi}_{1,t},\dots,\bm{\pi}_{m,t}]^{T},\bm{\nu}_{t}\coloneqq[\nu_{1,t},\dots,\nu_{m,t}]^{T}$ denote the auxiliary states and control inputs of system (16). We can define a sequence of functions as follows:
@@
where $i\in\{0,...,m-1\}.$ Since $a_{i,t}$ is a DHOCBF with relative degree m+1im+1-i for (16), based on (19), we define a constraint set $\mathcal{U}_{\bm{a}}$ for $\bm{\nu}_{t}$ as follows:
@@
Let $\psi_{i}(\mathbf{x},\bm{\Pi}),\ i\in\{1,...,m\}$ be defined by (22) and $\mathcal{C}_{i},\ i\in\{0,...,m-1\}$ be defined by (23). A function $h(\mathbf{x}):\mathbb{R}^{n}\to\mathbb{R}$ is a Discrete-Time Auxiliary-Variable Adaptive Control Barrier Function (DAVCBF) with relative degree $m$ for system (1) if every $a_{i,t},i\in\{1,...,m\}$ is a DHOCBF with relative degree m+1im+1-i for the auxiliary system (16), and there exist class $\kappa$ functions $\alpha_{j},j\in\{1,...,m\}$ and Lipschitz controller $\mathbf{u}_{t}$ s.t.
@@ IV-B Risk Sensitive Forward Invariance
$\rho_{0:t}(0,0,\dots,\psi_{i}(\mathbf{x}_{t},\bm{\Pi}_{t}))\geq 0,\quad\forall t\in\mathbb{N}_{\geq 0}.$ (26)
@@ IV-C Ensuring Risk Sensitive Safety with RACBFs
Let $\psi_{i}(\mathbf{x},\bm{\Pi}),\ i\in\{1,...,m\}$ be defined by (22), $\mathcal{C}_{i},\ i\in\{0,...,m-1\}$ be defined by (23) and a dynamic coherent risk measure $\rho$ be defined by Def. 4. A function $h(\mathbf{x}):\mathbb{R}^{n}\to\mathbb{R}$ is a Risk-Aware Adaptive Control Barrier Function (RACBF) with relative degree $m$ for system (13) if every $a_{i,t},i\in\{1,...,m\}$ is a DHOCBF with relative degree m+1im+1-i for the auxiliary system (16), and there exist class $\kappa$ functions $\alpha_{j},j\in\{1,...,m\}$ and Lipschitz controller $\mathbf{u}_{t}$ for $\psi_{i}(\mathbf{x},\bm{\Pi})$ s.t.
@@ IV-D Optimal Control with RACBFs
$\begin{split}-(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})\geq 0,\sigma_{i,t}\geq 0,\\ \sigma_{i,t}\geq-\rho^{i}(\psi_{m}(\mathbf{x}_{t+1},\bm{\Pi}_{t+1}))-\zeta_{t},i\in\{1,...,|\mathcal{W}|\}.\end{split}$ (31)
@@
where $\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})=\frac{\psi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{i,t}},\alpha_{i}(\psi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t}))=\lambda_{i}a_{i,t}\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t}),0<\lambda_{i}\leq 1,i\in\{1,\dots,m\}$. The term $\frac{a_{i,t}}{a_{i,t+1}}$ can be adjusted to ameliorate the conservativeness of control strategy that $(1-\lambda_{i})\phi_{i-1}(\mathbf{x}_{t},\bm{\Pi}_{t})$ may have, which allows for a broader range of feasible control inputs and confirms the adaptivity of DAVCBFs to control constraint and conservativeness of control strategy. The aforementioned benefits can be further amplified by formulating the optimal control problem in the model predictive control framework.
@@
Compared to Eqn. (32), we observe that $(1-\lambda_{i}a_{i,t})$ has a range of $\infty,1]$, while $\frac{a_{i,t}}{a_{i,t+1}}(1-\lambda_{i})\phi_{i-1}$ ranges from $0,+\infty]$. The form of Eqn. (32) maximizes the flexibility in adjusting the parameters of the class $\kappa$ function while eliminating negative parameters that may compromise safety.
@@ V CASE STUDY AND SIMULATIONS
where $M$ denotes the mass, and $\mathbf{x}_{t}=[x_{t},y_{t},\theta_{t},v_{t}]^{T}$ captures the 2-D location, heading angle, and linear speed; $\mathbf{u}_{t}=[u_{1,t},u_{2,t}]^{T}$ represents angular velocity ($u_{1,t}$) and driven force ($u_{2,t}$), respectively. The disturbance $\mathbf{w}_{t}\in\mathcal{W}$ enters the system linearly and is characterized by a probability mass function over the states. In this study, the probability mass function is a simple normal distribution centered at 0 with a standard deviation of $[0.2,0.2,0.2,0.2]$ for the four states $x_{t},y_{t},\theta_{t},v_{t}]$. The system is discretized with $\Delta t=0.1s$ and the total number of discretization steps $t_{T}$ in Problem. 1 is 50. System (34) is subject to the following input constraint:
@@
The initial state is $1.5,0,\frac{\pi}{12},1]^{T}$ and the target state is $\mathbf{x}_{e}=[2,0,0,0]^{T},$ which are marked as black diamond and green circle in Fig. 1. The circular obstacle is centered at $(x_{o},y_{o})=(0,0)$ with $r_{o}=1$, which is displayed in pale red. To initialize the proposed RACBFs, we multiply an auxiliary variable $a_{1,t}$ by a quadratic distance function for circular obstacle avoidance as $\psi_{0}(\mathbf{x}_{t},\bm{{\pi}}_{1,t})=a_{1,t}((x_{t}-x_{o})^{2}+(y_{t}-y_{o})^{2}-r_{o}^{2})$. Since the relative degree of $\psi_{0}$ with respect to system (34) is 2, the DHOCBFs for $a_{1,t}$ are defined as